OpenStudy (anonymous):

verify this identity (1-sin x)/(1+sin x)=(sec x-tan x)^2

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

i did i got stuck

OpenStudy (anonymous):

i am guessing multiply the left hand side top and bottom by \(1-\sin(x)\) might be snappy, but maybe not let me try it

OpenStudy (anonymous):

yeah that will meet @Loser66 in the middle if you multiply \[\frac{1-\sin(x)}{1+\sin(x)}\times \frac{1-\sin(x)}{1-\sin(x)}\] you get \[\frac{1-2\sin(x)+\sin^2(x)}{\cos^2(x)}\]

OpenStudy (anonymous):

yes thats where i got to and i got stuck

OpenStudy (anonymous):

oh isnt that what you get when you expand the right hand side?

OpenStudy (anonymous):

i am sure it is check it

OpenStudy (anonymous):

nope the rhs is sec ^2 x-2secx +tanx+tan^2 x

OpenStudy (anonymous):

\[\sec^2(x)-2\tan(x)\sec(x)+\tan^2(x)\]

OpenStudy (anonymous):

your plus should be a times in the middle

OpenStudy (anonymous):

that is what you get if you divide term by term from here \[\frac{1-2\sin(x)+\sin^2(x)}{\cos^2(x)}\]

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

get with the latex brother alph

OpenStudy (anonymous):

OMG @primeralph he got you dude....

OpenStudy (primeralph):

@satellite73 Why, why you always got to give away my secrets?